# Evolutionary Algorithms Simulating Molecular Evolution: A New Field Proposal

### v1

@@ 2 Abstract
The genetic blueprint for the essential functions of life is encoded in DNA, which is translated into proteins — the engines driving most of our metabolic processes. Recent advancements in genome sequencing have unveiled a vast diversity of protein families, but compared to the massive search space of all possible amino acid sequences, the set of known functional families is minimal. One could say nature has a limited protein “vocabulary.” The major question for computational biologists, therefore, is whether this vocabulary can be expanded to include useful proteins that went extinct long ago, or maybe never evolved in the first place. We outline a computational approach to solving this problem. By merging evolutionary algorithms, machine learning (ML), and bioinformatics, we can facilitate the development of completely novel proteins which have never existed before. We envision this work forming a new sub-field of computational evolution we dub evolutionary algorithms simulating molecular evolution (EASME).
@@ 3 Antecedents of EASME
Evolutionary algorithms (EAs), as the name suggests, model the natural process of evolution: with selection acting on variation where fitter individuals, i.e., individuals better adapted to their environment, have a higher chance of reproducing and surviving, thus increasing the odds of their genes proliferating within the population’s gene pool. Natural sexual reproduction capitalizes on recombination between individuals and stochastic variation to generate new alleles, thus diversifying the gene pool over many generations. By mimicking this process in silico, EAs have proven to be powerful engineering problem solvers, outperforming human experts on many optimization and design problems, frequently through unintuitive – or even counterintutitive – solution approaches. In 2004, Wolfgang Banzhaf and a team of eight other experts in the fields of artificial evolution and evolutionary biology, proposed a bold new idea – that progress in the field had slowed, but that it could continue to improve by incorporating more of the nuances of the natural world:
@@
Closing this loop at the molecular level is our goal by encoding individuals in EASME projects in molecular (DNA) terms, with fitness evaluations analyzing functional capabilities of those encoded proteins through de-novo folding and bioinformatic algorithms.
@@ 4 Defining the Problem
Proteomics — the study of proteins — is a primary focus point for the EASME field. Protein strings are sentences written with an alphabet of 20 amino acids. Many of these amino acid strings exceed 1,000 characters in length, thus the search space of possible protein strings is unfathomably vast. Most combinations of amino acids would do absolutely nothing and would be unstable. Thus the search space of possible protein configurations is a vast “sea of invalidity.” Within that sea exists a tiny archipelago of functional proteins, and only a small region of one of those islands is occupied by the proteins that actually evolved and remain extant today (see 1). EASME aims to expand the set of extant proteins by colonizing new islands in the sea of invalidity, yielding functional protein strings that can later be produced and analyzed in a lab.
@@ 7 Evolutionary Algorithms Simulating Molecular Evolution
As computing power continues to increase, complex simulations of evolving biochemical systems are approaching feasibility, even on desktop computers. This has been demonstrated most recently in “Modelling the emergence of Wolbachia Toxin antidote protein functions with an evolutionary algorithm” [2], a project in which researchers drove forward evolution of two specific, interacting proteins involved in cytoplasmic incompatibility, and observed the speed, location, and addition of novel functions that can be mapped to phenotypes. In another article, “Molecular dynamics simulation of an entire cell” by Stevens et al., an entire minimal cell was simulated in silico (stopping just short of full simulation of metabolic activity, which still proved difficult) [7]. With the computing power to simulate complex biochemical systems now at our disposal, and EAs that model the same process that produces those biomolecules, the stage is now set for the EASME field to bloom.
Through the continued development of this burgeoning field — leveraging the biomimicry of EAs (in particular genetic programming), ML, and state-of-the-art hardware — we feel uncovering grammar structures that define the fundamental languages of our biology is now possible (see 2). While this work will undoubtedly be complex, with processes like alternative splicing complicating the rules of any complete protein grammar, we are confident that utilizing all the AI tools now at our disposal will nonetheless make it feasible.
@@
Figure 2: EASME exists at the intersection of computational evolution, computational biology, and molecular evolution, drawing on the strengths of all three to solve complex problems.
@@ 8 Building and Testing the EASME Algorithm
Our hypothesis is that an EASME model would be composed of several algorithmic components. At the base of EASME would be an EA that would instantiate diverse populations of DNA-encoded proteins. The base EA would evolve and mutate genes the same way they naturally evolve — by point mutations, deletions, insertions, and recombinations. Fitness of encoded proteins would be determined by a multifaceted fitness function analyzing protein schemas (that bioinformatically define enzymatic consensus sequences) and protein grammar rules that would structurally validate a protein’s primary sequence based on de novo folding algorithms minimizing a free energy function and/or primary string attribute properties like hydrophobicity, isoelectric charge, amino acid sub-words, et cetera. Provided that a basic truthful understanding of protein grammar could be encoded, all evolved proteins in the population could be driven to evolve new functional clades, and eventually protein families. The fitness function could also be coupled with a “protein spam filter” that could rule out structurally unstable permutations and combinations, effectively reducing the search space of garbage that needs to be explored.
To reduce the complexity of designing the initial EASME, the algorithm could drive evolution of very important select proteins, such as key photosynthetic enzymes (for applied use in agriculture); but as EASME matures, any proteins could be evolved and optimized, and all functional islands in the sea of invalidity might be colonized over time. The final problem EASME will face is understanding how to determine the function of newly evolved proteins derived from in silico processes. Our suggestion is that libraries of EASME-derived peptides would first be chemically synthesized, then screened for useful activities. For example, a library of EASME proteins might first be screened against insects, to find EASME-derived proteins capable of killing insects. The positive hits could then be developed into novel insecticides for agriculture, which would also reinforce and improve the grammar structures of the fitness functions. Thus, we outline a feasible path toward the development of the EASME algorithm.
@@ 9 EASME to Date
While evolutionary methods have been applied to the fields of biology before, EASME represents a specific effort to focus on expanding the set of extant proteins. One could visualize evolutionary computation and biology intersecting along two axes — abstraction of means (representation, mutation operators, etc.), and abstraction of ends (nature of evolved organisms, etc.). For example, traditional genetic programming has a clear abstraction of means, with the evolved genes consisting of computer operations rather than nucleic acids. Genetic programming also represents an abstraction of ends, as genetic programming (typically) produces computer programs, not biological programs (genes, proteins, et cetera). The aforementioned cytoplasmic incompatibility research, in contrast, represents the distinction of EASME — both means and ends were represented in biological terms, rather than abstract terms. The individuals being evolved were actual codon sequences, evaluated based on their homology to the real-world Wolbachia genome, in order to model their actual evolution. We aim to push the boundaries of genetic programming even further, and use this same methodology to solve problems that have thus far been infeasible to solve in silico. At the moment, biological programs (proteins) cannot be efficiently “executed” on a computer, but we believe the same process that made them efficiently executable in the real world could make them efficiently executable in the digital world.
@@
10 Limitless Potential: Other Use Cases for EASME
@@
One of EASME’s grand challenges is the accurate prediction of actual protein functionality. While projects like AlphaFold have brought protein structure prediction closer to reality (albeit with the shortcomings we describe in Section 5), functionality is another matter entirely. EASME, however, may hold the potential to truly solve this problem. A project spearheaded by Wolfgang Banzhaf – the “father” of computational evolution, as described in Section 3 – has recently produced supporting evidence. As detailed in their paper “A Genetic Programming Approach to Engineering MRI Reporter Genes,” Banzhaf and his team have been able to use a new protein optimization tool, POET, to design 27 short peptides that produce more contrast in an MRI than was previously state-of-the-art [3]. POET utilizes genetic programming to achieve these results, and while this project could still be considered somewhat narrow in scope (evolving peptides towards one particular function rather than predicting the function of a completely arbitrary peptide), it still demonstrates that evolutionary methods hold potential for solving this problem, and the same methodology could easily be transferred to numerous other functions.

### v2

@@ 2 Abstract
The genetic blueprint for the essential functions of life is encoded in DNA, which is translated into proteins — the engines driving most of our metabolic processes. Recent advancements in genome sequencing have unveiled a vast diversity of protein families, but compared to the massive search space of all possible amino acid sequences, the set of known functional families is minimal. One could say nature has a limited protein “vocabulary.” A major question for computational biologists, therefore, is whether this vocabulary can be expanded to include useful proteins that went extinct long ago, or have never evolved (yet). By merging evolutionary algorithms, machine learning, and bioinformatics, we can develop highly-customized “designer proteins.” We dub the new sub-field of computational evolution which employs evolutionary algorithms with DNA string representations, biologically-accurate molecular evolution, and bioinformatics-informed fitness functions, Evolutionary Algorithms Simulating Molecular Evolution (EASME).
@@ 3 Antecedents of EASME
Evolutionary algorithms (EAs), as the name suggests, model the natural process of evolution: with selection acting on variation where fitter individuals, i.e., individuals better adapted to their environment, have a higher chance of reproducing and surviving, thus increasing the odds of their genes proliferating within the population’s gene pool. Natural sexual reproduction capitalizes on recombination between individuals and stochastic mutation to generate new alleles, thus diversifying the gene pool over many generations. By mimicking this process in silico, EAs have proven to be powerful engineering problem solvers, outperforming human experts on many optimization and design problems, frequently through unintuitive – or even counterintutitive – solution approaches. In 2006, Wolfgang Banzhaf and a team of eight other experts in the fields of artificial evolution and evolutionary biology, proposed a bold new idea – that progress in the field had slowed, but that it could continue to improve by incorporating more of the nuances of the natural world:
@@
Closing this loop at the molecular level is our goal by simulating DNA evolution in EASME projects, with fitness evaluations analyzing functional capabilities of those encoded proteins through de-novo folding and bioinformatic algorithms.
@@ 4 Defining the Problem
Proteomics — the study of proteins — is a primary focus point for the EASME field. Protein strings are sentences written with an alphabet of 20 amino acids. Many of these amino acid strings exceed 1,000 characters in length, thus the search space of possible protein strings is unfathomably vast. Most combinations of amino acids would do absolutely nothing and would be unstable. Thus the search space of possible protein configurations is a vast “sea of invalidity.” Within that sea exists a tiny archipelago of functional proteins, and only a small region of one of those islands is occupied by the proteins that actually evolved and remain extant today (see Figure 1). EASME aims to expand the set of extant proteins by colonizing new islands in the sea of invalidity, yielding functional protein strings that can later be produced and analyzed in a lab.
@@ 7 Evolutionary Algorithms Simulating Molecular Evolution
As computing power continues to increase, complex simulations of evolving biochemical systems are approaching feasibility, even on desktop computers. This has been demonstrated most recently in “Modeling the emergence of Wolbachia toxin antidote protein functions with an evolutionary algorithm” [2], a project in which researchers drove forward evolution of two specific, interacting proteins involved in cytoplasmic incompatibility, and observed the speed, location, and addition of novel functions that can be mapped to phenotypes. In another article, “Molecular dynamics simulation of an entire cell” by Stevens et al., an entire minimal cell was simulated in silico (stopping just short of full simulation of metabolic activity, which still proved difficult) [7]. With the computing power to simulate complex biochemical systems now at our disposal, and EAs that model the same process that produces those biomolecules, the stage is now set for the EASME field to bloom.
Through the continued development of this burgeoning field — leveraging the biomimicry of EAs (in particular genetic programming), ML, and state-of-the-art hardware — we feel uncovering grammar structures that define the fundamental languages of our biology is now possible (see Figure 2). While this work will undoubtedly be complex, we are confident that utilizing all the AI tools now at our disposal will nonetheless make it feasible.
@@
Figure 2: EASME exists at the intersection of computational evolution, computational biology, and molecular evolution, drawing on the strengths of all three to solve problems.
@@ 8 Building and Testing the EASME Algorithm
Our hypothesis is that an EASME model would be composed of several algorithmic components. At the base of EASME would be an EA that would instantiate diverse populations of DNA sequences. The base EA would evolve and mutate genes the same way they naturally evolve — by point mutations, deletions, insertions, and recombinations. Fitness of encoded proteins would be determined by a multifaceted fitness function analyzing protein schemas (that bioinformatically define enzymatic consensus sequences) and protein grammar rules that would structurally validate a protein’s primary sequence based on de novo folding algorithms minimizing a free energy function and/or primary string attribute properties like hydrophobicity, isoelectric charge, amino acid sub-words, et cetera. Provided that a basic truthful understanding of protein grammar could be encoded, all evolved proteins in the population could be driven to evolve new functional clades, and eventually protein families. The fitness function could also be coupled with a “protein spam filter” that could rule out structurally unstable permutations and combinations, effectively reducing the search space of garbage that needs to be explored (see Figure 3).
Figure 3: Intelligent application of ML can help constrain the search space of proteins.
To reduce the complexity of designing the initial EASME, the algorithm could drive evolution of very important select proteins, such as key photosynthetic enzymes (for applied use in agriculture); but as EASME matures, any proteins could be evolved and optimized, and all functional islands in the sea of invalidity might be colonized over time.
The final problem EASME will face is understanding how to determine the function of newly evolved proteins derived from in silico processes. Our suggestion is that libraries of EASME-derived peptides would first be chemically synthesized, then screened for useful activities. For example, a library of EASME proteins might first be screened against insects, to find EASME-derived proteins capable of killing insects. The positive hits could then be developed into novel insecticides for agriculture, which would also reinforce and improve the grammar structures of the fitness functions. Thus, we outline a feasible path toward the development of the EASME algorithm.
@@ 9 EASME to Date
While evolutionary methods have been applied to the fields of biology before, EASME represents a specific effort to focus on biologically-accurate representations, with one compelling application being expanding the set of extant proteins. Most EAs applied to biology were abstract top-down population dynamic models. To date, EASME is the first to attempt modeling of actual DNA sequences. For example, the aforementioned cytoplasmic incompatibility research represents an initial development of EASME where both means and ends were represented in biological terms, rather than abstract terms. The individuals being evolved were actual codon sequences, evaluated based on their homology to the real-world Wolbachia genome, in order to model their actual evolution. We aim to push the boundaries of genetic programming even further, and use this same methodology to solve problems that have thus far been infeasible to solve in silico. While molecular cellular simulations of complete organisms are still computationally intractable, simulating their evolution is feasible.
@@
10 The Path Forward: Other Use Cases for EASME
Figure 4: Ways to employ EASME.
EASME represents a specific effort to focus on expanding the set of extant proteins. The EASME framework, once developed, can run in two distinct ways (see Figure 4). First, EASME can evolve a random sequence toward a known consensus sequence (“unknown to known”). In this context, the desired outcome is to reconstruct sequence clusters that went extinct during the process of evolution. Selective fitness is implemented by pushing the evolution towards a known protein sequence family. EASME outputs samples of Pareto optimal sequences from theoretical evolutionary intermediates, effectively recovering extinct sequence variants. How much the EASME generated sequences would differ from real historical intermediates is unknowable without ancient genomes. However, the utility of generated sequences can be tested, measured, and linked to a corresponding successful discovery rate. The second way to run EASME is known to unknown, where a known entity is forward evolved into the future by implementing a selection regimen that drives towards a desired characteristic phenotype. This methodology outputs Pareto optimal sequences that may have never evolved yet and is effectively a fast forward button on evolution into the future. While this approach would undoubtedly produce many false positives, wet lab work will allow us to test and validate designed proteins while simultaneously honing a given enzyme’s fitness function. Biologically measuring the ratio of valid to invalid protein outputs would allow us to optimize the design process (and even if that ratio is low, it will still be orders of magnitude faster than natural evolution, a process which plays out on evolutionary timescales). To achieve both these ends, EASME will employ EA and GP models supplemented with ML where appropriate. As an aside, EASME can also be run in reverse, where genes and genomes currently extant are reverse-engineered in attempt to reconstruct ancient common ancestors.
@@ 10 The Path Forward: Other Use Cases for EASME
One of EASME’s grand challenges is the accurate prediction of protein functionality. While projects like AlphaFold have brought protein structure prediction closer to reality (albeit with the shortcomings we describe in Section 5), functionality is another matter entirely. EASME, however, may hold the potential to truly solve this problem. A project spearheaded by Wolfgang Banzhaf – the “father” of computational evolution, as described in Section 3 – has recently produced supporting evidence. As detailed in their paper “A Genetic Programming Approach to Engineering MRI Reporter Genes,” Banzhaf and his team have been able to use a new protein optimization tool, POET, to design 27 short peptides that produce more contrast in an MRI than was previously state-of-the-art [3]. POET utilizes genetic programming to achieve these results, and while this project could still be considered somewhat narrow in scope (evolving peptides towards one particular function rather than predicting the function of a completely arbitrary peptide), it still demonstrates that evolutionary methods hold potential for solving this problem, and the same methodology could easily be transferred to numerous other functions.
@@ 11 What Needs to Happen Next?
We plan to document the progress of the EASME field at https://aub.ie/easme. We welcome contributions from any researchers with an interest in this field.
Acknowledgements
The authors gratefully acknowledge Wolfgang Banzhaf’s insightful feedback on this paper.